# Radiant Triangle Soup with Soft Connectivity Forces for 3D Reconstruction and Novel View Synthesis

### v1

@@ 7 Limitations & Conclusions
Our geometric loss terms, much like 2DGS, involves a trade-off between model reconstruction quality and novel-view synthesis. Adhering more strictly to the geometric losses in our framework can lead to over-smoothing in some areas, and at times can prohibit the movement of primitives, as triangles are more strongly attracted to each other, leaving parts of the scene under-reconstructed. Furthermore, due to the periodic nearest-neighbors search, the run-time for our method increases with the number of primitives. Because of this, it can become difficult to reconstruct very large scenes.
In this work, we introduce a new scene representation, namely Radiant Triangle Soup (RTS). To the best of our knowledge, we are the first to introduce explicit 3D forces between primitives in a splatting framework, helping to coordinate the positioning of primitives to directly form surfaces. The introduction of Triangle Soup as the scene representation for radiance fields is amenable to future work in surface optimization and can be augmented to include more extensive forms of triangle connectivity and mesh evolution. After some modifications, our framework can also be used for optimization over watertight meshes. In future work, we plan to extend RTS with modified primitive connectivity strategies and mesh post-processing refinement.
Supplementary Material
S.1 Implementation Details
In this section, we describe our experimental setting and optimization parameters in detail. To begin optimization, similar to previous work [4, 2], all geometric supervision is disabled, with optimization only being guided initially by the SSIM loss $\mathcal{L}_{ssim}$. We enable the normal consistency loss $\mathcal{L}_{norm}$ at iteration $7,000$ and enable both the smoothness loss $\mathcal{L}_{smooth}$ and connectivity loss $\mathcal{L}_{conn}$ at iteration $10,000$, both chosen empirically. All triangles start with an initial opacity ($\alpha$) value set to $0.1$, with opacity for all primitives being reset every $3,000$ iterations. We run optimization on the scenes from the DTU and Mip-NeRF 360 datasets for $25,000$ and $30,000$ iterations, respectively. Densification and pruning is run every $250$ iterations starting after iteration $2,000$. The maximum reference point gradient threshold for densification is set to $7.5e^{-5}$ in all experiments.
@@
  • Reference Point ($\mu$): $\begin{bmatrix}1.5e^{-4},&2e^{-6}\end{bmatrix}$
@@
where the Reference Point learning rate follows an exponential decay scheduler starting with $1.5e^{-4}$ and ending with $2e^{-6}$ conditioned on the total number of iterations. We also tested a linear decay scheduler leading to similar results.
@@
S.2 Ablation on Depth Rendering
We show an ablation on the effects of using mean and median depth, in Table S.1. While using median depth has virtually no effect on the quality of the reconstructed images, there is a substantial effect on the accuracy and consistency of the rendered depth maps, and thus, the final geometry of the 3D model. The normal consistency loss $\mathcal{L}_{norm}$ is indifferent to surface layering when computed w.r.t. the blended mean depth and the blended mean normal. Using the median depth guides the median triangle to better align with the blended normals and vice versa.
Table S.1: Comparison of the effects of mean vs median depth using the entire DTU evaluation set.
Method Acc.(mm) $\downarrow$ Comp.(mm) $\downarrow$ PSNR $\uparrow$ Primitives $\downarrow$
Mean Depth 1.23 1.36 31.69 228,000
Median Depth 0.64 0.70 31.68 249,000
Table S.2: Novel View Synthesis on all scenes from the Mip-NeRF 360 [42] dataset.
Method Outdoor Indoor
@@
NeRF [11] 21.46 0.458 0.515 26.84 0.790 0.370
Deep Blending [43] 21.54 0.524 0.364 26.40 0.844 0.261
Instant NGP [16] 22.90 0.566 0.371 29.15 0.880 0.216
MipNeRF360 [42] 24.47 0.691 0.283 31.72 0.917 0.180
SuGaR [3] 22.93 0.629 0.356 29.43 0.906 0.225
3DGS [1] 24.64 0.731 0.234 30.41 0.920 0.189
2DGS [2] 24.34 0.717 0.246 30.40 0.916 0.195
GOF [13] 24.82 0.750 0.202 30.79 0.924 0.184
PGSR [4] 24.76 0.752 0.203 30.36 0.934 0.147
3DCS [5] 24.07 0.700 0.238 31.33 0.927 0.166
RTS 21.41 0.657 0.349 30.28 0.921 0.130
S.3 Additional Evaluations
In this subsection, we show additional results on the Mip-NeRF 360 dataset [42]. Following the protocol specified by the authors of Mip-NeRF 360, we separate the images in each scene, taking every eighth image as a test image and training on the remaining. As standard evaluation, we report PSNR, SSIM, and LPIPS [41] metrics. We show results on all indoor and outdoor scenes in Table S.2. For indoor scenes, RTS shows results on par with the state-of-the-art methods, having the leading LPIPS score, the fourth overall SSIM score, and a highly competitive PSNR score among all listed methods. On two outdoor scenes (treehill and flowers), RTS shows difficulty reconstructing extremely distant background foliage, impacting the overall metrics.
We provide supplemental renderings of novel views on challenging indoor and outdoor scenes with fine details and non-Lambertian surfaces. Output renderings and 3D models for all scenes will be made publicly available to the research community if the paper is accepted.
We show qualitative geometric reconstruction results on all scenes from the DTU dataset in Fig. S.2 and novel view synthesis results on all indoor and outdoor scenes from the Mip-NeRF 360 datasets in Fig. S.1 and Fig. S.3, respectively.
@@
Figure S.1: Visualizations of all indoor scenes, (top-down) [bonsai, counter, kitchen, room], from the Mip-NeRF 360 dataset [42].
Figure S.2: Point cloud visualization for all scenes from the evaluation set from the DTU dataset [6].
@@
Figure S.3: Visualizations of all outdoor scenes, (top-down) [bicycle, flowers, garden, stump, treehill], from the Mip-NeRF 360 dataset [42].

### v2

@@ 6 Limitations & Conclusions
In this work, we introduce a new scene representation, namely Radiant Triangle Soup (RTS). To the best of our knowledge, we are the first to introduce explicit 3D forces between primitives in a splatting framework, helping to coordinate the positioning of primitives to directly form surfaces. Modifying the weights of these forces allows for tuning between 3D reconstruction quality and novel-view synthesis quality. The main limitation of our current algorithm is its inability to extract watertight meshes. Furthermore, due to the periodic nearest-neighbors search, there is a minor increase in run-time proportional to the number of primitives. Please see Section S.2 for more details.
The introduction of Triangle Soup as the underlying representation for Radiance Fields is amenable to future work in surface optimization. We plan to extend RTS with modified primitive connectivity strategies and perform optimization over watertight meshes.
Supplemental Material
Here we include additional material on background geometry used for computing barycentric coordinates, further implementation details, and additional results.
S.1 Background
In this section, we introduce the necessary background geometry used to parameterize a ray-triangle intersection in barycentric coordinates. The barycentric coordinates of a point on a triangle $\lambda=[\lambda^{0},\lambda^{1},\lambda^{2}]^{T}$, s.t. $\sum_{j=1}^{3}\lambda^{j}=1$, provide a means for expressing the coordinates of the point as a linear combination of the coordinates of the three vertices, $V^{0},V^{1},V^{2}]$. This parameterization is important for graphics applications to be able to efficiently rasterize triangles onto screen space and interpolate the color of a pixel from each vertex.
In rendering, we compute the ray-triangle intersections using,
$P=C+(\widehat{r}d)$ (S.1)
$d=\frac{\widehat{n}\cdot\overrightarrow{CB}}{\widehat{n}\cdot\widehat{r}}$ (S.2)
$\lambda=\frac{1}{\widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}V_{2}})}\begin{bmatrix}\widehat{n}\cdot(\overrightarrow{V_{1}V_{2}}\times\overrightarrow{V_{1}P})\\ \widehat{n}\cdot(\overrightarrow{V_{2}V_{0}}\times\overrightarrow{V_{2}P})\\ \widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}P})\end{bmatrix}$ (S.3)
S.2 Implementation Details
In this section, we describe our experimental setting and optimization parameters in detail. To begin optimization, similar to previous work (Huang et al., 2024; Chen et al., 2024a), all geometric supervision is disabled, with optimization only being guided initially by the SSIM loss $\mathcal{L}_{ssim}$. We enable the normal consistency loss $\mathcal{L}_{norm}$ at iteration $7,000$ and enable both the smoothness loss $\mathcal{L}_{smooth}$ and connectivity loss $\mathcal{L}_{conn}$ at iteration $10,000$, both chosen empirically. All triangles start with an initial opacity ($\alpha$) value set to $0.1$, with opacity for all primitives being reset every $3,000$ iterations. We run optimization on the scenes from the DTU and Mip-NeRF 360 datasets for $25,000$ and $30,000$ iterations, respectively. Densification and pruning is run every $250$ iterations starting after iteration $2,000$. The maximum incenter gradient threshold for densification is set to $7.5e^{-5}$ in all experiments.
@@
  • Incenter ($\mu$): $\begin{bmatrix}1.5e^{-4},&2e^{-6}\end{bmatrix}$
@@
where the incenter learning rate follows an exponential decay scheduler starting with $1.5e^{-4}$ and ending with $2e^{-6}$ conditioned on the total number of iterations. We also tested a linear decay scheduler leading to similar results.
Edge Connectivity Overhead
To compute neighboring edge connections, we construct a single KD-Tree containing the midpoints of all triangle edges which is then queried for each edge once. The connected edge indices are the only structure that is stored during optimization. Since the parameters of the primitives are modified during optimization, we need to recompute the KD-Tree and connected edge indices every 250 iterations (aligned with Adaptive Density Control), which ultimately leads to a minimal overhead for moderate size scenes. To be concrete, building the KD-Tree and computing the neighboring indices takes on average 4 seconds for around 300,000 primitives, which is more than the average number of primitives needed to reconstruct the scenes for the DTU dataset. Since this operation only happens every 250 iterations starting after iteration 10,000 (when the connectivity loss is activated), the overhead of this operation only adds roughly 5 minutes to the optimization process, which is about a $6\%$ increase in runtime. Reconstructing larger scenes, or more precisely, scenes that require more primitives, will naturally demand a larger overhead.
Run-time
@@
Figure S.1: Point cloud visualization for all scenes from the evaluation set from the DTU dataset(Aanæs et al., 2016).
S.3 Loss Functions
For completeness, we define the loss terms used in this paper that were introduced in previous work (Kerbl et al., 2023; Huang et al., 2024). The SSIM loss is computed as follows:
$\mathcal{L}_{ssim}=\frac{1}{N}\sum_{i,j}(1-\gamma)L_{1}+\gamma L_{D\text{-}SSIM}$ (S.4)
The normal consistency loss is computed as follows:
$\mathcal{L}_{norm}=\frac{1}{N}\sum_{i,j}(1-\widehat{n}^{T}\widehat{n}_{d})$ (S.5)
This encourages alignment between the rendered normal $\widehat{n}$ with the normal computed from the rendered depth map $\widehat{n}_{d}$,
S.4 Additional Evaluations
Table S.1: Novel View Synthesis per-scene on all scenes from the mip-NeRF 360 dataset (Barron et al., 2022).
Scene SSIM PSNR LPIPS
room 0.919 31.25 0.156
counter 0.900 28.27 0.139
kitchen 0.936 30.69 0.083
bonsai 0.930 30.93 0.142
bicycle 0.637 20.31 0.365
flowers 0.617 19.99 0.351
garden 0.742 24.89 0.250
stump 0.658 23.25 0.410
treehill 0.631 18.64 0.371
In Table S.1, we show the per-scene novel view synthesis evaluation for RTS on the mip-NeRF 360 dataset (Barron et al., 2022) using the standard metrics of PSNR, SSIM, and LPIPS. As von Lützow & Nießner (2025) note, using primitives with explicit boundaries can begin to introduce hard edges in regions with poor visibility while smoother primitives degrade more gracefully. While the diffuse boundaries of our triangles help prevent much of this behavior, its effects are noticeable in some of the reconstructed images, especially in outdoor scenes with distant background foliage. See Fig. S.3
As mentioned in Section 5.3, we compare the Chamfer distances of RTS and two competitive GS algorithms in Table S.2 with and without the use of the GT segmentation masks. While the distances increase for all methods, this experiment demonstrates how RTS is more effective at floater removal and background modeling.
We provide additional renderings of novel views on challenging indoor and outdoor scenes with fine details and non-Lambertian surfaces. We show qualitative geometric reconstruction results on all scenes from the DTU dataset in Fig. S.1 and novel view synthesis results on all indoor and outdoor scenes from the Mip-NeRF 360 datasets in Fig. S.2 and Fig. S.3, respectively.
Table S.2: Chamfer distance evaluation on scenes from DTU (Aanæs et al., 2016). Following previous literature, we average the accuracy and completeness Chamfer distances on the widely used evaluation set. Chamfer distances are measured in millimeters. Top: Chamfer distance evaluation using the GT mask. Bottom: Chamfer distance evaluation without using the GT mask.
Method DTU
@@
2DGS (Huang et al., 2024) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
PGSR (Chen et al., 2024a) 0.36 0.57 0.38 0.33 0.78 0.58 0.50 1.08 0.63 0.59 0.46 0.54 0.30 0.38 0.34 0.52
RTS 0.42 0.61 0.74 0.39 0.53 0.86 0.70 0.84 0.72 0.37 0.68 0.87 0.34 0.58 0.44 0.61
2DGS (Huang et al., 2024) [no mask] 1.22 1.69 0.88 0.43 0.96 0.77 0.85 1.23 1.87 1.69 0.91 2.01 0.82 0.82 1.07 1.15
PGSR (Chen et al., 2024a) [no mask] 1.10 1.59 0.97 0.45 1.98 0.64 0.59 1.70 1.69 1.57 0.74 0.63 0.44 0.71 0.74 1.04
RTS [no mask] 0.52 0.92 0.94 0.55 0.91 1.01 0.74 1.09 1.05 0.67 0.88 1.07 0.42 0.79 0.80 0.82
@@
Figure S.2: Visualizations of all indoor scenes, (top-down) [bonsai, counter, kitchen, room], from the mip-NeRF 360 dataset (Barron et al., 2022).
@@
Figure S.3: Visualizations of all outdoor scenes, (top-down) [bicycle, flowers, garden, stump, treehill], from the Mip-NeRF 360 dataset (Barron et al., 2022). In the top-right section of the rendered image of treehill, RTS imprecisely approximates the appearance of the background foliage, while GS-based algorithms typically blur this region.